THEORETICAL CALCULATION OF THE TORSIONAL SPECTRUM OF THE PARTIALLY DEUTER-ATED SPECIES OF METHANOL

S. BLASCO, D. LAUVERGNAT, Laboratoire de Chimie Physique, Unité mixte du C.N.R.S. et de l’Université Paris-Sud, Bât. 490, Université Paris-Sud, 91405 Orsay Cedex, France; AND \underline{L. H. COUDERT}, Laboratoire de Photophysique Moléculaire, C.N.R.S., Bât. 350, Université Paris-Sud, 91405 Orsay Cedex, France.

The first spectroscopic investigations of the partially deuterated species of methanol, CH_{2}DOH and CD_{2}HOH, have shown that these species display a dense torsional spectrum^{a} difficult to assign. With a view toward understanding these spectra, a theoretical calculation of their rotation-torsion energy levels has been undertaken aided by ab initio calculations. This calculation accounts for the complicated torsion-rotation interaction displayed by these molecules and is based on the following features:

• The kinetic energy part of the Hamiltonian is calculated numerically taking into account all 12 internal degrees of freedom of the molecules. The angle of internal rotation is treated as an active coordinate.^{b}

• The Schrödinger equation for the internal rotation is solved using Gaussian quadrature.^{c}

• Internal degrees of freedom corresponding to the other small amplitude motions are treated using the harmonic approximation, for each value of the internal angle of rotation corresponding to the DVR grid.^{d}

• The potential energy function of the molecule is obtained using ab initio calculations. After making some reasonable assumptions for the dipole moment function, the last step of the calculation involves evaluating the intensity of the rotation-torsion transitions. In the paper we hope to be able to show plots of roomtemperature absorption spectra.

aMukhopadhyay, Perry, Butler, Herbst, and De Lucia, 57th International Symposium on Molecular Spectroscopy, paper RH06 (2002) and Mukhopadhyay, Perry, Lock, and Klee, 57th International Symposium on Molecular Spectroscopy, paper RH07 (2002).

bLauvergnat and Nauts, J. Chem. Phys. 116, 8560 (2002).

cLight and Bačić, J. Chem. Phys. 87, 408 (1987).

dLauvergnat, Nauts, Justum and Chapuisat, J. Chem. Phys. 114, 6592 (2001).